OpenStudy (anonymous):

Can somebody please help me with this! (Will give medals) m/4+m/4=5/8

OpenStudy (mathstudent55):

You can easily add the left side together since both fractions have the same denominator.

OpenStudy (mathstudent55):

Then cross multiply like you'd do with a proportion.

OpenStudy (anonymous):

Umm. i did something terribly wrong sorry it's m/4+m/2=5/8

OpenStudy (bmorse):

2m/4=5/8...multiply both sides by 4 2m=20/8 now divide both sides by 2 m= 20/16 = 5/4

OpenStudy (anonymous):

I still kind of don't understand.

OpenStudy (anonymous):

I am going to write this on paper.

OpenStudy (bmorse):

We want a common denominator m/4+m/2=5/8 so all the denominators should be 8

OpenStudy (anonymous):

Oh ok i starting to understand.

OpenStudy (bmorse):

2m/8+4m/8 = 5/8 now you can add the common fractions.. 6m/8=5/8 multiply both sides by 8 6m=5 m=5/6

OpenStudy (bmorse):

by multiplying both sides by 8 you get rid of the denominators

OpenStudy (anonymous):

I am starting to get it, i am just writing it on paper.

OpenStudy (bmorse):

yes write it out so you can see it better and let me know if you have any questions or get stuck

OpenStudy (anonymous):

Ok i will let you know. Thank you!